# Optimal Algorithm for Multiple-Processor Multitask Scheduling

### v1

@@ 3 $\mathsf{NP}$-$\mathsf{SRPT}$ Algorithm and Competitive Ratio Analysis
$r^{\mathsf{NP}-{\pi}}(d,a)$
[…]
@@
$r^{\mathsf{NP-SRPT}}(p,a)$
[…]
@@
3.1 A general upper bound on relevant work for $\mathsf{NP}$-$\pi$ policy
@@
$z^{\pi}_{\mathrm{J}}(y,t)=[\sup\{a\in[0,p_{\mathrm{J}}]|r^{\pi}(p_{\mathrm{J}},a)\leq y\}-a_{\mathrm{J}}(t)]\cdot\mathbbm{1}({r^{\pi}_{\mathsf{J}}\leq y}).$
@@
$\mathsf{Rel}^{\pi}_{\leq y}(t)=\sum_{\mathrm{J}}{z^{\pi}_{\mathrm{J}}(y,t)}.$
@@
$z^{\pi}_{y}=\sup\{a\geq 0|r^{\pi}(p_{\mathrm{J}},p_{\mathrm{J}}-a)\leq y\}.$
Here we extend the definition of old job age cutoff in [24] for allowing monotone decreasing policy. It can be seen that $z^{\pi}_{\mathrm{J}}(y,t)\leq z^{\pi}_{y}$ holds for all monotone policy $\pi$.
@@
$\mathsf{Rel}^{\mathsf{NP}-\pi}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)\leq(N-1)\cdot(z^{\pi}_{y}+\eta),\forall r,t\geq 0,$
@@
Proof: We analyze the following two cases:
  • few-jobs time intervals, in which there are $N-1$ or fewer jobs of rank no more than $y$. Consider one of such job $\mathsf{J}$, if one of the non-preemptive tasks of this job is under processing, then
    $z^{\mathsf{NP}-\pi}_{\mathrm{J}}(y,t)\leq z^{\pi}_{y}+\eta,$
@@
    $z^{\mathsf{NP}-\pi}_{\mathrm{J}}(y,t)\leq\eta.$
    If this job does not have a non-preemptive task under processing, then we have $z^{\mathsf{NP}-\pi}_{\mathrm{J}}(y,t)\leq z^{\pi}_{y}$.
@@
    $\mathsf{Rel}^{\mathsf{NP}-\pi}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)$
    $\leq$ $\mathsf{Rel}^{\mathsf{NP}-\pi}_{\leq y}(t)\leq(N-1)\cdot(z^{\pi}_{y}+\eta).$
  • many-jobs time intervals, in which all $N$ servers are serving jobs of rank no more than $y$. Note that all $N$ servers are working towards reducing the value of $\mathsf{Rel}^{\mathsf{NP}-\pi}_{\leq y}(t)$, while the two systems experience the same arrival sequence. Therefore, in the many interval period, the rate of decrease of $\mathsf{Rel}^{\mathsf{NP}-\pi}_{\leq y}(t)$ in the $N$-server system is no less than the rate of decrease of $\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)$ in single server system. Consequently, we have the following:
    $\mathsf{Rel}^{\mathsf{NP}-\pi}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)$
    $\leq$ $\mathsf{Rel}^{\mathsf{NP}-\pi}_{\leq y}(t^{{\dagger}})-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t^{{\dagger}})\leq(N-1)\cdot(z^{\pi}_{y}+\eta),$
@@
$\Box$
3.2 Performance Analysis of $\mathsf{NP}$-$\mathsf{SRPT}$
@@
$\mathsf{W}^{\mathsf{NP-SRPT}}_{\leq y}(t)-\mathsf{W}^{\pi^{*}_{1,N}}_{\leq y}(t)\leq N\cdot(z^{\pi}_{y}+\eta).$
Proof: As SRPT is a monotone decreasing SOAP policy, we have
$z^{\mathrm{SRPT}}_{y}=y.$
@@
$\mathsf{Rel}^{\mathsf{NP-SRPT}}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)\leq(N-1)\cdot(z^{\pi}_{y}+\eta).$ (1)
@@
$\mathsf{Rel}_{\leq y}^{\pi}(t)=\mathsf{W}_{\leq y}^{\pi}(t)+\mathsf{V}_{\leq y}^{\pi}(t).$ (2)
@@
$\mathsf{W}^{\mathsf{NP-SRPT}}_{\leq y}(t)-\mathsf{W}^{\pi^{*}_{1,N}}_{\leq y}(t)$
$=$ $\mathsf{Rel}^{\mathsf{NP-SRPT}}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)+\mathsf{V}^{\pi^{*}_{1,N}}_{\leq y}(t)-\mathsf{V}^{\mathsf{NP-SRPT}}_{\leq y}(t)$
$\leq$ $\mathsf{Rel}^{\mathsf{NP-SRPT}}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)+\mathsf{V}^{\pi^{*}_{1,N}}_{\leq y}(t).$
Since under policy $\pi^{*}_{1,N}$, there is at most one job with a rank of 00, and its relevant $y$-work does not exceed $\eta+y$. Combined with (1), the proof is complete. $\Box$
@@
$n_{\mathsf{NP-SRPT}}(t)\leq n_{\pi^{*}_{1,N}}(t)+N\cdot(\ln\alpha+\beta+1).$
Proof: In Appendix A we present a self-contained proof, which shows that $n_{\mathsf{NP-SRPT}}(t)\leq n_{\pi^{*}_{1,N}}(t)+4N\cdot(\log\alpha+\beta+1)$. The main idea is to divide the jobs into different classes and compare the remaining number of jobs under $\mathsf{NP}$-$\mathsf{SRPT}$ with that under optimal algorithm $\pi^{*}_{1,N}$. For any algorithm $\pi$, at time slot $t$, we divide the unfinished jobs into $\Theta(\log\alpha)$ classes $\{\mathcal{C}_{k}(\pi,t)\}_{k\in[\log\alpha+1]}$, based on their remaining workload. Jobs with remaining workload that is no more than $2^{k}$ and larger than $2^{k-1}$ are assigned to the $k$-th class. Formally,
@@
holds for $G/G/N$ system.
$n_{\mathsf{NP-SRPT}}(t)$ $\leq\int_{0}^{\infty}{\frac{\mathbbm{E}[\mathsf{Rel}^{\pi^{*}}_{\leq r}]}{r^{2}}}\mathrm{d}r+\int_{0}^{\infty}{\frac{N\max\{r,\eta\}}{r^{2}}}\mathrm{d}r$
@@
$\mathcal{CR}_{\mathsf{NP-SRPT}}\leq\ln\alpha+\beta+1.$
Proof: Again we divide the time slots into few-job intervals and many-job intervals, the competitive ratio of $\mathsf{NP}$-$\mathsf{SRPT}$ satisfies that
$\mathcal{CR}_{\mathsf{NP-SRPT}}$
$=$ $\frac{\int_{t:n_{\mathsf{NP-SRPT}}(t)<N}{n_{\mathsf{NP-SRPT}}(t)}}{F^{\pi^{*}_{1,N}}}$
$+\frac{\int_{t:n_{\mathsf{NP-SRPT}}(t)\geq N}{n_{\mathsf{NP-SRPT}}(t)}}{F^{\pi^{*}_{1,N}}}$
$\leq$ $\frac{\int_{t:n_{\mathsf{NP-SRPT}}(t)\geq N}{n_{\pi^{*}}(t)}}{F^{\pi^{*}_{1,N}}}+\Big{(}\ln\alpha+\beta\Big{)}\cdot\frac{\int_{t:n_{\mathsf{NP-SRPT}}(t)\geq N}{N}}{F^{\pi^{*}_{1,N}}}$
$+\frac{\int_{t:n_{\mathsf{NP-SRPT}}(t)<N}{n_{\mathsf{NP-SRPT}}(t)}}{F^{\pi^{*}_{1,N}}}$
@@
where the first inequality follows from Lemma 10, the second inequality is due to the fact that $\int_{t:n_{\mathsf{NP-SRPT}}(t)\geq N}{N}+\int_{t:n_{\mathsf{NP-SRPT}}(t)<N}{n_{\mathsf{NP-SRPT}}(t)}$ and $\int_{t:n_{\mathsf{NP-SRPT}}(t)\geq N}{n_{\pi^{*}_{1,N}}(t)}$ are two lower bounds of minimum total response time. As $F^{\pi^{*}_{1,N}}\leq F^{\pi^{*}}$, the proof is complete. $\Box$
3.3 Competitive ratio lower bound
@@
Proof: When $\eta=0$, the problem degenerates to preemptive setting and no algorithm can achieve a competitive ratio of $o(\ln\alpha)$ [25]. When $\eta=p_{\max}$, the problem degenerates to the non-preemptive setting and $O(\beta)$ is the best possible competitive ratio if the number of machines is constant [28]. The proof is complete. $\Box$
Proposition 14.
@@
Proof: The reasoning is similar as the proof of Proposition 13, since work-conserving algorithms cannot achieve a competitive ratio of $o(\beta^{1-\varepsilon})$ in the non-preemptive single-task job scheduling [28]. $\Box$
4 Asymptotic Optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ with Poisson Arrival
@@
Let $F^{\mathsf{NP-SRPT}}_{\rho}$ and $F^{\pi^{*}}_{\rho}$ be the response time incurred by $\mathsf{NP}$-$\mathsf{SRPT}$ and optimal algorithm respectively, when the traffic intensity is equal to $\rho$. In an M/G/N with finite task size and job size distribution satisfying $\mathbbm{E}[p^{2}(\log p)^{+}]<\infty$, $\mathsf{NP}$-$\mathsf{SRPT}$ is heavy traffic optimal, i.e.,
$\lim\nolimits_{\rho\rightarrow 1}\frac{\mathbbm{E}[F^{\mathsf{NP-SRPT}}_{\rho}]}{\mathbbm{E}[F^{\pi^{*}}_{\rho}]}=1.$ (3)
@@
4.1 Average response time bound
Our important step is to derive the following analytical upper bound on $\mathbbm{E}[F_{\rho}^{\mathsf{NP-SRPT}}]$.

### v2

@@ III $\mathsf{NP}$-$\mathsf{SRPT}$ Algorithm and Competitive Ratio Analysis
$r^{\mathsf{NP}\text{-}{\pi}}(d,a)$
$=$ $r^{\pi}(d,a)\cdot\mathbbm{1}(\nexists\text{ started but unfinished non-preemptive task}).$
@@
$r^{\mathsf{NP\text{-}SRPT}}(p,a)$
$=$ $(p-a)\cdot\mathbbm{1}(\nexists\text{ started but unfinished non-preemptive task}),$
@@
III-A A general upper bound on relevant work for $\mathsf{NP}$-$\pi$ policy
@@
$z^{\pi}_{\mathsf{J}}(y,t)$
$=$ $[\sup\{a\in[0,p_{\mathsf{J}}]|r^{\pi}(p_{\mathsf{J}},a)\leq y\}-a_{\mathsf{J}}(t)]\cdot\mathbbm{1}({r^{\pi}_{\mathsf{J}}(p_{\mathsf{J}},a_{\mathsf{J}}(t))\leq y}),$
where $a_{\mathsf{J}}(t)$ denotes the age of job $\mathsf{J}$ at time $t$.
@@
$\mathsf{Rel}^{\pi}_{\leq y}(t)=\sum_{\mathsf{J}}{z^{\pi}_{\mathsf{J}}(y,t)}.$
@@
$z^{\pi}_{\mathsf{J}}(y)=\sup\{a\geq 0|r^{\pi}(p_{\mathsf{J}},p_{\mathsf{J}}-a)\leq y\}.$
Here we extend the definition of old job age cutoff in [38] for allowing monotone decreasing policy. It can be seen that
$z^{\pi}_{\mathsf{J}}(y,t)\leq z^{\pi}_{\mathsf{J}}(y)$
holds for all monotone policy $\pi$.
@@
$\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)\leq(N-1)\cdot(\max_{\mathsf{J}}z^{\pi}_{\mathsf{J}}({y})+\eta),\forall r,t\geq 0,$
@@
Proof.
We analyze the following two cases:
  • few-jobs time intervals, in which there are N1N-1 or fewer jobs of rank no more than $y$. Consider one of such job $\mathsf{J}$, if one of the non-preemptive tasks of this job is under processing, then
    $z^{\mathsf{NP}\text{-}\pi}_{\mathsf{J}}(y,t)\leq\max_{\mathsf{J}}z^{\pi}_{\mathsf{J}}({y})+\eta,$
@@
    $z^{\mathsf{NP}\text{-}\pi}_{\mathsf{J}}(y,t)\leq\eta.$
    If this job does not have a non-preemptive task under processing, then we have $z^{\mathsf{NP}\text{-}\pi}_{\mathsf{J}}(y,t)\leq\max_{\mathsf{J}}z^{\pi}_{\mathsf{J}}({y})$.
@@
    $\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)$
    $\leq$ $\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq y}(t)\leq(N-1)\cdot(\max_{\mathsf{J}}z^{\pi}_{\mathsf{J}}({y})+\eta).$
  • many-jobs time intervals, in which all $N$ servers are serving jobs of rank no more than $y$. Note that all $N$ servers are working towards reducing the value of $\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq y}(t)$, while the two systems experience the same arrival sequence. Therefore, in the many interval period, the rate of decrease of $\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq y}(t)$ in the $N$-server system is no less than the rate of decrease of $\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)$ in single server system. Consequently, we have the following:
    $\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)$
    $\leq$ $\mathsf{Rel}^{\mathsf{NP}\text{-}\pi}_{\leq y}(t^{{\dagger}})-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t^{{\dagger}})$
    $\leq$ $(N-1)\cdot(\max_{\mathsf{J}}z^{\pi}_{\mathsf{J}}({y})+\eta),$
@@
III-B Performance Analysis of $\mathsf{NP}$-$\mathsf{SRPT}$
@@
$\mathsf{W}^{\mathsf{NP\text{-}SRPT}}_{\leq y}(t)-\mathsf{W}^{\pi^{*}_{1,N}}_{\leq y}(t)\leq N\cdot(y+\eta).$
Proof.
As SRPT is a monotone decreasing SOAP policy, we have
$z^{\mathrm{SRPT}}_{\mathsf{J}}(y)=y,\forall\mathsf{J}.$
@@
$\mathsf{Rel}^{\mathsf{NP\text{-}SRPT}}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)\leq(N-1)\cdot(y+\eta).$ (1)
@@
$\mathsf{Rel}_{\leq y}^{\mathsf{NP}\text{-}\mathsf{SRPT}}(t)=\mathsf{W}_{\leq y}^{\mathsf{NP}\text{-}\mathsf{SRPT}}(t)+\mathsf{V}_{\leq y}^{\mathsf{NP}\text{-}\mathsf{SRPT}}(t).$ (2)
@@
$\mathsf{W}^{\mathsf{NP\text{-}SRPT}}_{\leq y}(t)-\mathsf{W}^{\pi^{*}_{1,N}}_{\leq y}(t)$
$=$ $\mathsf{Rel}^{\mathsf{NP\text{-}SRPT}}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)+\mathsf{V}^{\pi^{*}_{1,N}}_{\leq y}(t)-\mathsf{V}^{\mathsf{NP\text{-}SRPT}}_{\leq y}(t)$
$\leq$ $\mathsf{Rel}^{\mathsf{NP\text{-}SRPT}}_{\leq y}(t)-\mathsf{Rel}^{\pi^{*}_{1,N}}_{\leq y}(t)+\mathsf{V}^{\pi^{*}_{1,N}}_{\leq y}(t).$
Since under policy $\pi^{*}_{1,N}$, there is at most one job with a rank of 0, and its relevant $y$-work does not exceed $\eta+y$. Combined with (1), the proof is complete. ∎
@@
$n_{\mathsf{NP\text{-}SRPT}}(t)\leq n_{\pi^{*}_{1,N}}(t)+N\cdot(\ln\alpha+\beta+1).$
Proof.
In Appendix Proof of Lemma 3 we present a self-contained proof, which shows that $n_{\mathsf{NP\text{-}SRPT}}(t)\leq n_{\pi^{*}_{1,N}}(t)+N\cdot(\ln\alpha+\beta+1)$. The main idea is to divide the jobs into different classes and compare the remaining number of jobs under $\mathsf{NP}$-$\mathsf{SRPT}$ with that under optimal algorithm $\pi^{*}_{1,N}$. For any algorithm $\pi$, at time slot $t$, we divide the unfinished jobs into $\Theta(\log\alpha)$ classes $\{\mathcal{C}_{k}(\pi,t)\}_{k\in[\log\alpha+1]}$, based on their remaining workload. Jobs with remaining workload that is no more than $2^{k}$ and larger than $2^{k-1}$ are assigned to the $k$-th class. Formally,
@@
holds for $\mathsf{G}/\mathsf{G}/\mathsf{N}$ system.
$n_{\mathsf{NP\text{-}SRPT}}(t)$ $\leq\int_{0}^{\infty}{\frac{\mathbbm{E}[\mathsf{Rel}^{\pi^{*}}_{\leq r}]}{r^{2}}}\mathrm{d}r+\int_{0}^{\infty}{\frac{N(r+\eta)}{r^{2}}}\mathrm{d}r$
@@
$\mathcal{CR}_{\mathsf{NP\text{-}SRPT}}\leq\ln\alpha+\beta+1.$
Proof.
Again we divide the time slots into few-job intervals and many-job intervals, the competitive ratio of $\mathsf{NP}$-$\mathsf{SRPT}$ satisfies that
$\mathcal{CR}_{\mathsf{NP\text{-}SRPT}}$
$=$ $\frac{\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)<N}{n_{\mathsf{NP\text{-}SRPT}}(t)}}{F^{\pi^{*}_{1,N}}}$
$+\frac{\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)\geq N}{n_{\mathsf{NP\text{-}SRPT}}(t)}}{F^{\pi^{*}_{1,N}}}$
$\leq$ $\frac{\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)\geq N}{n_{\pi^{*}}(t)}}{F^{\pi^{*}_{1,N}}}+\Big(\ln\alpha+\beta\Big)\cdot\frac{\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)\geq N}{N}}{F^{\pi^{*}_{1,N}}}$
$+\frac{\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)<N}{n_{\mathsf{NP\text{-}SRPT}}(t)}}{F^{\pi^{*}_{1,N}}}$
@@
where the first inequality follows from Lemma 3, the second inequality is due to the fact that
$\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)\geq N}{N}+\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)<N}{n_{\mathsf{NP\text{-}SRPT}}(t)}$
and
$\int_{t:n_{\mathsf{NP\text{-}SRPT}}(t)\geq N}{n_{\pi^{*}_{1,N}}(t)}$
are two lower bounds of minimum total response time. As $F^{\pi^{*}_{1,N}}\leq F^{\pi^{*}}$, the proof is complete. ∎
III-C Competitive ratio lower bound
@@
Proof.
When $\eta=0$, the problem degenerates to preemptive setting and no algorithm can achieve a competitive ratio of $o(\ln\alpha)$ [24]. When $\eta=p_{\max}$, the problem degenerates to the non-preemptive setting and $O(\beta)$ is the best possible competitive ratio if the number of machines is constant [8]. The proof is complete. ∎
Fact 2.
@@
Proof.
The reasoning is similar as the proof of Proposition 1, since work-conserving algorithms cannot achieve a competitive ratio of $o(\beta^{1-\varepsilon})$ in the non-preemptive single-task job scheduling [8]. ∎
IV Asymptotic Optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ with Poisson Arrival
@@
Let $F^{\mathsf{NP\text{-}SRPT}}_{\rho}$ and $F^{\pi^{*}}_{\rho}$ be the response time incurred by $\mathsf{NP}$-$\mathsf{SRPT}$ and optimal algorithm respectively, when the traffic intensity is equal to $\rho$. In an $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$ with finite task size and job size distribution satisfying $\mathbbm{E}[p^{2}(\log p)^{+}]<\infty$, $\mathsf{NP}$-$\mathsf{SRPT}$ is heavy traffic optimal, i.e.,
$\lim\nolimits_{\rho\rightarrow 1}\frac{\mathbbm{E}[F^{\mathsf{NP\text{-}SRPT}}_{\rho}]}{\mathbbm{E}[F^{\pi^{*}}_{\rho}]}=1.$ (3)
@@
IV-A Average response time bound
Our important step is to derive the following analytical upper bound on $\mathbbm{E}[F_{\rho}^{\mathsf{NP\text{-}SRPT}}]$.